A room is 1.5 times as long as it is wide, and its perimeter is 75 feet. Find the dimensions of the room.
Answer by checkley77(12844)

L=1.5W
2L+2W=P
2*1.5W+2W=75
3W+2W=75
5W=75
W=75/5
W=15 FT. IS THE WIDTH.
L=1.5*15
L=22.5 FT. ANS. FOR THE LENGTH.
PROOF:
2*22.5+2*15=75
45+30=75
75=75